### v2

# A Mathematical Definition of Path Integrals on Symplectic Manifolds

Abstract
We give a mathematical definition of some path integrals, emphasizing those relevant to the quantization of symplectic manifolds (and more generally, Poisson manifolds) — in particular, the coherent state path integral. We show that Kähler manifolds provide many computable examples and we emphasize those whose Bergman kernel is constant along the diagonal.
Contents
- 0 Introduction
1 The Simplest Case
- 1.1 Definition of the Path Integral
- 1.2 The Hilbert space and Path Integral – Operator Correspondence
- 1.3 Example: Conventional Quantum Mechanics
- 1.4 Relation to Deformation Quantization and Special Observables
- 1.5 Non-Uniqueness of Primitives
- 1.6 The Derivative of the Measure on Paths
- 1.7 Extending the Fundamental Theorem of Calculus to Path Measures
- 2 The General Case: The Propagator of Symplectic Manifolds
- 3 A New Look at the Riemann Integral
- 4 Examples
- 5 Relation to Other Quantization Schemes
0 Introduction
We discuss a mathematical definition of some path integrals, particularly those relevant to the quantization of symplectic manifolds such as
for In eq. 0.0.1, is the connection on the prequantum line bundle and the integrand should be interpreted as parallel transport over from the vector space over to the vector space over 222Implicit in this notation is a choice of orthonormal bases of the vector spaces over We have written it this way to be consistent with the physics literature, eg. equation (2.7) of [11]. See section 2 for the definition. Computing this path integral immediately produces a quantization, ie. a good Hilbert space and quantization map defined on all classical observables, as explained in section 1, section 2. We will call it a propagator.
We suggest that computing eq. 0.0.1 is equivalent to finding a propagator satisfying definition 2.0.1. In eq. 0.0.2, the integral is over maps from the disk into where are points on the boundary of the disk. It is related to eq. 0.0.1, since
Equation 0.0.2 naturally generalizes to Poisson manifolds. Kähler manifolds produce many computable examples, with particularly nice ones arising from those whose Bergman kernel is constant along the diagonal. As such this is closely related to Toeplitz quantization [3], as discussed in [6], [28].
Equation 0.0.1 is sometimes called the coherent state path integral (or phase space path integral) and Daubechies, Klauder in [7], [16] computed it on surfaces of constant curvature using Brownian motion — see section 4.2, where we also discuss the issue of uniqueness.333It was more generally computed on Kähler manifolds in [6] — when we say it was computed, we mean a reasonable value was assigned to it. A general definition isn’t given. In some sense it is the most natural and mathematically well-behaved form of quantization, and as we will discuss, it is closely related to the different geometric and deformation quantization programs, [35], [36], [19], [10]. In particular, eq. 0.0.2 is closely related to the Poisson sigma model, see [5] remark 3, and [2] equation (3.39). Also see [12].
The idea is that we can compute eq. 0.0.1 by finding a function on with the right derivative. We will not prove any very general existence theorems, but we do obtain well-defined existence and uniqueness questions. This paper is largely a continuation of [22], where we formalized the procedure of putting path integrals over morphisms of Lie algebroids on a lattice (eg. path integrals over maps between manifolds, section 4.2.2). That paper describes a path integral construction analogous to the construction of the Riemann integral via Riemann sums, eg. Feynman’s construction of the path integral. This paper, on the other hand, focuses on the path integral analogue of the fundamental theorem of calculus — the construction of the propagator fits into the framework of the aforementioned paper.
In particular, we have the following lemma:
Lemma.
Indeed, the examples computed by Daubechies, Klauder all have the property that the Bergman kernel is constant along the diagonal. The significance in geometric quantization of it being constant has been known at least since [4] (where they also give a sufficient condition for it to be constant), and such prequantizations were studied by Donaldson in [8], where they are called balanced. In general, the normalized Bergman kernel integrates the prequantum connection of the Bergman Kähler form, which is an -perturbation of the original Kähler form. It is a rescaling of the pullback of the Fubini-Study form by the Kodaira embedding, [26], [33].
Idea:
The definition we give is simple, and to understand it we will begin with an analogy with the definition of antiderivatives, which will lead us into eq. 1.1.2, 1.1.3:
Let be a 1-form on An antiderivative is normally said to be a function such that However, an antiderivative can be equivalently defined as a function
which satisfies the following two conditions:
The fundamental theorem of calculus says that
Suppose that we want to find such an antiderivative. In order to compute it, we can first choose any function satisfying eq. 0.0.6, ignoring eq. 0.0.5. We can then triangulate the interval with vertices and consider the sum
This gives an approximation to whose limit over subdivisions converges to The traditional choices for are the ones giving the left-point and right–point Riemann sums:
However, any choice of smooth satisfying eq. 0.0.6 will do, see proposition 3.1.1. This gives a notion of Riemann sum which behaves well with respect to pullbacks, which is needed for the path integral. We can think of a path integral as an antiderivative of a differential form in a different category — a category of complex measures. Feynman’s way of computing them is the analogue of taking a limit of Riemann sums.
Note that, we can replace the additive structure used in eq. 0.0.5 with a multiplicative structure:
However, this doesn’t really give anything different, but we do get something different if we only ask that eq. 0.0.10 holds “on average”. This leads us to the path integral.
Remark 0.0.1.
Stated differently, an antiderivative of is a 1-cocycle on the pair groupoid
such that where is the van Est map. See appendix .3.
1 The Simplest Case
1.1 Definition of the Path Integral
Let be a manifold with a measure and consider a purely imaginary 1-form Let
be a smooth function, considered as an integral kernel with respect to Suppose that satisfies the following three conditions (compare with eq. 0.0.5, 0.0.6): for all
The first condition says that the convolution and this is one of the defining properties of a (time-independent) propagator;444See remark 1.1.2 for a comment on the time-dependent case. is the exterior derivative in the second factor; the additional third and fourth conditions are normalization and reversal of orientation properties.555These condition is automatically satisfied by where satisfies eq. 0.0.5 and with the multiplicative structure replaced by the additive structure.
The first property can be thought of as a cocycle condition on the pair groupoid and the second condition says that where is the van Est map, see appendix .3 (note that, is non-zero on a neighborhood of the identity bisection).
Definition 1.1.1.
If satisfies the previous four conditions, then we say that integrates (as a propagator), or that differentiates to
should be thought of as the amplitude of a particle beginning at to end at 666With a zero Hamiltonian. To understand in what sense is a path integral, suppose that
Using eq. 1.1.2, we have that
hence
Now pick a smooth path and pick a triangulation of Due to eq. 1.1.3 (see proposition 3.1.1), we have that
Hence, taking in eq. 1.1.8, we can formally say that
Now, if we ignore condition eq. 1.1.2, then the limit of the right side of eq. 1.1.8 is essentially Feynman’s method of defining eq. 1.1.10, which is made precise by the lattice framework developped in [22], see section 4.2.2. Formally then, the right side of eq. 1.1.8 is a Riemann sum with the algebraic property of eq. 0.0.5 replaced by eq. 1.1.2 (of course, we have to understand Riemann sums more generally to make this statement, see section 3).
By the complex version of Komolgorov’s extension theorem (A.2 of [25]), such an defines a complex measure on the space of paths though a priori this measure is concentrated on all paths (including discontinuous ones).777One should be able to modify the measurable space so that this measure is concentrated on a smaller set of discontinuous paths. A relevant discussion is in Berezin [1]. In the case of a non-trivial line bundle, as in the next section, the measure is operator-valued. We can denote this measure by
Most of the observables in quantum theory are given by correlation functions and are of the form
for some functions and This is equal to
Remark 1.1.2.
In the time-dependent case, eg. for Brownian motion or a non-zero Hamiltonian, eq. 1.1.2 is replaced by
1.2 The Hilbert space and Path Integral – Operator Correspondence
This section puts the work of [16] into general context. We can use a non-trivial line bundle with Hermitian form rather than the trivial one, and we will do this in the next section.
From we get the following sesquilinear form999We assume linearity in the second component. on
which we assume is bounded (this is automatic if ). The Riesz representation theorem gives an equivalence between bounded sesquilinear forms and bounded linear operators on Hilbert spaces, and the corresponding operator is given by
That is, given a sesquilinear form on a Hilbert space there is a unique linear operator such that
The conditions assumed on imply that is an orthogonal projection, ie. and from which it follows that
Definition 1.2.1.
The space of physical states (or quantum states), denoted is the eigenspace of
On eq. 1.2.1 agrees with the -inner product. This eigenspace is analogous to a space of Kähler-polarized sections of a prequantum line bundle, and we can identify it with where if
Such are sometimes called reproducing kernels and have been widely studied, eg. [15], [29]. has the special property that pointwise evaluation is a bounded linear functional, since by the definition of for
for all Therefore, by the Riesz representation theorem, has “delta functions” for each point denoted That is, if then there are genuine states such that
Definition 1.2.2.
The states are the coherent states
Coherent states form an overcomplete basis of in the sense that they are linearly dependent.
It follows from the definition of eq. 1.2.1 and the fact that for each that
where is the dual linear functional of
Definition 1.2.3.
There is a quantization map
given by associating to each the following sesquilinear form:
This is the “matrix element” of the observable determined by the path integral (for any ) and is given by
Explicitly,
To be clear, the correspondence between the path integral and the operator formulations is an instance of the correspondence determined by the Riesz representation theorem between sesquilinear forms and operators.101010However, there is a sense in which the path integral formulation logically precedes the operator formalism. This is also one reason why the coherent space path integral (or phase space path integral, [18]) is nice, as this correspondence is obfuscated when a polarization is used.
These quantum operators have the property that hence
where on the left is the inner product of On we can write diagonally in the coherent state basis, ie.
Lemma 1.2.4.
From this, it follows that forms an irreducible representation of the quantum operators, so indeed is a good Hilbert space.
1.3 Example: Conventional Quantum Mechanics
Consider the symplectic manifold with coordinates and symplectic form Let
is a propagator integrating according to definition 1.1.1. This is equal to
The state is identified with the corresponding quantum mechanical coherent state, ie. the eigenvector of the lowering operator111111These are usually defined in the context of the harmonic oscillator. for which the expectation values of the momentum and position operators are respectively. The physical Hilbert space is the Segal-Bargmann representation (sometimes called the holomorphic representation), [17].
Many other examples come from Kähler manifolds, as we will discuss in section 4.2.
1.4 Relation to Deformation Quantization and Special Observables
1.4.1 Special Observables
There is a special class of functions for which
where the latter is equal to
Such functions generalize the harmonic functions on phase space (in particular, the coordinate functions ). There is a sense in which these observables are the most canonically quantizable ones — their classical and quantum expectation values agree on all states, in the sense of -algebras ([38]). That is, given a normalized radon measure on (ie. a mixed classical state), we get a mixed quantum state given by
and if satisfies eq. 1.4.1 then
On the left is a quantum expectation value and on the right is a classical expectation value.121212These describe states with classical statistical uncertainty. In [7], the dual quantization scheme is used, ie. a function quantizes to an operator such that However, such a quantization map doesn’t need to exist.
In special cases our quantization map
is an isomorphism of vector spaces, and in this case we get a noncommutative product on In some other case, there may be a special subspace of functions for which the quantization map is injective and algebraically closed, in which case one gets a noncommutative product on this special subspace.
1.4.2 Deformation Quantization
We can get a quantization of a symplectic manifold by choosing a propagator integrating the connection on the prequantum line bundle, as discussed in section 2. One can often show that the quantization map is perturbatively an isomorphism. Here, we should consider propagators integrating
In this case we obtain a formal deformation quantization, see equation (2.10) of [11]. This is what occurs in Toeplitz quantization [3]. See section 5.1, [28] for more on this relation. To make sense of what it measn to be an isomorphism perturbatively, we give the following definition:
Definition 1.4.1.
(see [14]) Let be a Poisson manifold and let be a set containing as an accumulation point. For each let be a unital -algebra such that and let
be ∗- linear131313This compatibility with ∗ can be relaxed. such that its image generates Furthermore, assume that is the inclusion map. We say that is a (non-perturbative) deformation quantization of if there is a star product on such that, for all
where is the component of the star product up to order
In this definition, we don’t have to assume that is a star product a priori, just that it is an asymptotic expansion in
1.5 Non-Uniqueness of Primitives
It is important to note that, while primitives in the sense of eq. 0.0.5 are unique (this uses connectivity of ), primitives in the sense of eq. 1.1.2 are not. Therefore, we are abusing notation by writing eq. 1.1.10 — we should really consider the right side to be the set of all functions satisfying the four aforementioned conditions and write
This is a similar abuse of notation one commits when stating that is the antiderivative of
We address the non-uniqueness more in the example of section 4.2, where we show that the different propagators are distinguished by their second order Taylor expansions, rather than the expected first order.
1.6 The Derivative of the Measure on Paths
Consider a complex measure on the space of (discontinuous) paths The condition corresponding to eq. 1.1.3 says that the derivative of the Radon-Nikodym derivative of the 2-dimensional distributions of equals — its 2-dimensional distribution being the complex measure on given by
That is, the integral above is over all paths such that
The Radon-Nikyodym derivative is taken with respect to the product measure This defines a function
and we require that
Of course, we are making some differentiability assumptions on the complex measure to do this. We generalize this to path integrals from higher dimensional simplices in section 3.3.
1.7 Extending the Fundamental Theorem of Calculus to Path Measures
The ultimate goal is to extend the fundamental theorem of calculus in this setting. The main issues to understand are:
1.
non-uniqueness of the propagators differentiating to (see section 4.2 for a discussion),
- 2.
The non-uniqueness is due to the fact that, under path integrals, we can’t really integrate differential forms. This is because to uniquely specify Riemann-type integrals under path integrals we need to specify higher order information, eg. classically we integrate objects such as but under path integrals we integrate objects such as
Here, can be considered as a symmetric tensor, or as a homogenous polynomial of degree 2. We could integrate such objects classically as well, but we don’t bother to make it explicit because the higher order terms don’t contribute to the integral. However, in path integrals they do contribute. Picking good higher order terms often seems necessary to get convergence. To be a bit more precise, for a partition of the interval classically
This is why the usual definition of the Riemann integral doesn’t depend on the choice of Riemann sum, ie. why the choice of in eq. 1.7.1 is never specified. However, eq. 1.7.2 isn’t quite true under path integrals and therefore we need to be more precise than just specifying a 1-form. We discuss this in detail in section 3.
2 The General Case: The Propagator of Symplectic Manifolds
More generally, we can replace with square-integrable sections of a complex line bundle with Hermitian connection, where of the previous section corresponds to the connection, and everything essentially works the same. Of course, the most natural setting for this is a prequantizable symplectic manifold which comes with a natural measure. Even better are Kähler manifolds, since there are canonically defined propagators, as discussed in section 4.
In this setting, consider a line bundle with Hermitian connection
In this case, the path integral we want to compute is
where denotes parallel transport between the vector spaces over and over the curve That is, eq. 2.0.2 must define a linear map between any two vector spaces of the line bundle. The kernel we need to consider then is a section of
where are the projections onto the first and second factor, respectively. Equation 2.0.3 defines a line bundle over and points in this line bundle are naturally identified with linear maps between the corresponding vector spaces of
The conditions of definition 1.1.1 are replaced by: for all
The subscript in eq. 2.0.5 means that the covariant derivative is being taken in the second component; each defines a section of given by
where the subscript is used to denote the corresponding vector space. Equation 2.0.5 says that the covariant derivative of this section at is zero.
Definition 2.0.1.
If satisfies the previous conditions and the corresponding sesquilinear form eq. 2.0.19 is bounded, we say that it integrates (as a propagator).151515We may relax eq. 2.0.4, see section 5.
Any section of eq. 2.0.3 which equals on the diagonal determines a connection on (by differentiating in the second component at the diagonal),161616We are using the identification of splittings with connections. and eq. 2.0.5 says that this connection is equal to Furthermore, for any smooth path and triangulation of it follows that
converges to parallel transport over as We discuss uniqueness of such a propagator in section 4.2.
Remark 2.0.2.
In the symplectic case eq. 0.0.1, we can interpret the exponent as specifying the connection only to leading order in in which case it is natural to relax eq. 2.0.5 to
See remark 4.1.3.
The next lemma tells us that that if we can find such an without assuming eq. 2.0.6, then we can always normalize so that all conditions are satisfied:
Lemma 2.0.3.
Suppose satisfies the following conditions: for all
Then is a propagator integrating where is given by the following: let then
The rest goes as in section 1, with the sesquilinear form eq. 1.2.1 replaced by
The only difference is that, due to nontriviality of the line bundle, the states are only defined up to a scalar. To define a coherent state, choose a normalized vector over We get a bounded linear functional on the physical Hilbert space which is given by square-integrable sections of the line bundle satisfying
where is the operator corresponding to the sesquilinear form eq. 2.0.19. Given such a section
for some The map
defines a bounded linear functional, by eq. 2.0.20. The Riesz representation theorem then defines the coherent states.
Definition 2.0.4.
The coherent state is defined to be the vector in corresponding to the linear functional eq. 2.0.22.
If then
Therefore, associated to there is a canonical projection operator, which we denote by 181818As pointed out to me by Francis Bischoff, the map is a version of the Kodaira embedding. In addition, it follows that
where is any vector over with unit norm.
Note that, it is only as a vector state that the coherent state isn’t well-defined, as the linear functional on operators (ie. a -algebra state)
it is well-defined. Of course, the associated projection operator is also well-defined as discussed above.
Interestingly, if we can find a subspace of sections of the prequantum line bundle for which pointwise evaluation is continuous, then we can try to flip this construction to determine a propagator via eq. 2.0.24. We will do this in section 4.1.
Remark 2.0.5.
Usually, a quantization is said to be something like an operator assignment on a Hilbert space , such that
(eg. [32]). However, it’s not clear that any such a structure should count as a quantization, since it appears to completely ignore any classical–quantum state correspondence. It is important that for any classical state 191919In the sense of -algebras, ie. is a continuous, positive linear functional on the -algebra. one can choose quantum states such that as More precisely, we should have that
Otherwise, one can’t recover classical dynamics, as Heisenberg’s equations only limit to Hamilton’s equations on such sequences of states. The coherent state path integral approach comes with such states.
3 A New Look at the Riemann Integral
3.1 The One-Dimensional Case
To make sense of our definition of the path integral, we need to consider a more mathematically natural version of the Riemann integral. In particular, we need a version which behaves well with respect to pullbacks, since the path integral involves pulling back forms over paths. To motivate the definition, consider the following:
Proposition 3.1.1.
Stated more geometrically, eq. 3.1.1 says that the exterior derivative of in the second factor, evaluated at the diagonal, is equal to
The usual choices of are These result in the left-point and right-point Riemmann sums, respectively. The left side of eq. 3.1.2 can be considered to be a generalized Riemann sum. In particular, if then
also satisfies the conditions, so this definition makes the fundamental theorem of calculus almost tautological. This notion of integration easily to give a notion of Riemann sums of differential forms on manifolds. This gives a notion of integration on manifolds which mirrors the notion of integration on coordinate space, ie. it involves triangulations and not partitions of unity or coordinates.
3.2 Integration Over Non-Differentiable Paths
Importantly, consider a smooth map This induces a map
and if satisfies the conditions of proposition 3.1.1 with respect to then satisfies the conditions of proposition 3.1.1 with respect to Therefore, if can be used to approximate
then can be used to approximate
As stated in section 1.5, while the integral of makes sense classically, it isn’t well-defined under path integrals because proposition 3.1.1 doesn’t hold there. This is related to the non-differentiable nature of paths in the path integral. The problem is that, different functions chosen in eq. 3.1.1 lead to different results under the path integral. Therefore, we need to specify more data. In the 1-dimensional Lagrangian formulation of the path integral (or better, its Euclidean version with respect to the Wiener measure), it is enough to specify
That is, we need to specify the second derivative of as well. This means that
is well-defined in Wiener space if for a path we take it to mean
where vanishes on the diagonal and
To state a precise result, we have the following well-known result:
Proposition 3.2.1.
212121The Riemann sum on the left defines the Stratonovich integral and the Riemann sum on the right defines the Itô integral, [31]. The former satisfies the fundamental theorem of calculus.With respect to the Wiener measure on continuous paths and for a smooth function
in
The left sum is the midpoint rule and the right sum is the left-point rule. The summands differ by a term of order Classically, this term doesn’t contribute to the integral because for smooth paths However, paths in Wiener space are generically Hölder continuous of exponent less than , and Since the paths are generically Hölder continous with exponent greater than ([9], [30]) it follows that all terms of order and higher are negligible.
Note that, even on a smooth path these left and right sums are not Riemann sums in the traditional sense. In order to be a Riemann sum, the summand would need to be multiplied by not However, in the more general sense of proposition 3.1.1, they are Riemann sums. The same issue arises in Feynman’s construction of the path integral. The sums involved in the approximations are not traditional Riemann sums, they are Riemann sums in our more general sense.
We will come back to this discussion in section 4.2.
Remark 3.2.2.
It would be interesting to know if there are forms which are integrable in this sense but which aren’t Lebesgue integrable. For example, consider a differentiable function such that isn’t Lebesgue integrable. It is still the case that
satisfies the conditions required to be used for a generalized Riemann sum of and that
converges in the limit (of course, to ). If we choose another function satisfying the conditions to be used for a Riemann sum of and such that
converges, must it converge to
3.3 Differentiating Path Measures to Differential Forms
Label the vertices of the (topological) standard -simplex by Suppose we have a complex measure on the space of maps with respect to some sigma algebra for which maps of the form
are measurable, for and for any Then we can consider the finite dimensional distributions given by
We can take the Radon-Nikodym derivative of these finite dimensional distributions with respect to and this gives us a function
Definition 3.3.1.
The derivative of is the -form on given by
If we write then takes the exterior derivative of in each of the components at See definition .3.1.
Of course, in the previous definition we are assuming that the finite dimensional distributions are absolutely continuous with respect to and that it is -times differentiable at the diagonal.
We suggest that any complex measure on the space of maps which formally equals
should have a derivative which is equal to
4 Examples
4.1 Toeplitz Quantization and Bergman Kernels
Let be a prequantizable Kähler manifold, with prequantum line bundle
Then pointwise evaluation of Kähler-polarized sections is a continuous map, and as previously discussed each unit-norm vector over determines a linear functional given by where
We denote by the associated section determined by the Riesz representation theorem. Note that, is independent of and we denote it by
We can define by
where have unit norm (this is independent of ).
The corresponding operator (section 1.2) is the orthogonal projection of onto the subspace of holomorphic sections. The quantization map determined by the path integral is equal to the Toeplitz quantization map [3]. Its corresponding integral kernel is given by
and this is polarized with respect to the Kähler polarization on 222222To be more clear, it is polarized with respect to a possibly different connection — it is this integral kernel multiplied by that is polarized with respect to (consistent with [13]).
is closely related to the Bergman kernel ([34]), which is important in the quantization of Kähler manifolds ([28]), and is called the Bergman measure (or coherent measure in [15]). The Bergman kernel can be described as follows: let be an orthonormal basis for the space of holomorphic sections. Then is a section of and is given by
We then define a propagator232323Note that, is canonically trivial over the diagonal, so we can identify with a positive-valued function.
Note that,
Lemma 4.1.1.
integrates with respect to the Bergman measure
if and only if is constant.
Proof.
We need to show that if and only if is constant on the diagonal. For the only if direction, differentiating along vectors in the eigenspace and using the product rule implies that is a holomorphic function. However, it is also real-valued, so it must be constant.
The other direction follows from product rule: since and is constant, for vectors we have that
Since is antiholomorphic in it follows that for in the eigenspace (ie. of the form Since it is also true for in the eigenspace (since is holomorphic in ) the result follows. ∎
Example 4.1.2.
is constant for where is the Fubini-Study symplectic form ([34]).
We get a function given by
This defines a function on composable arrows of the pair groupoid, and if is constant, then applying to eq. 5.1.2 gives This is closely related to deformation quantization. Equation 5.1.2 should be interpreted as
where the domain of integration is all maps with three marked points on mapping to (compare with equations (2.9), (2.11) of [11]). If is simply connected then
where is any map which restricts to on This corresponds to eq. 0.0.2 (after gauge reduction, see [2]).
For a relation of this quantization scheme to Berezin’s star product and the Poisson sigma model on Kähler manifolds, see [28].
Note that, these examples somewhat contradicts the discussion in section 2.3 of [11]. This is due to the fact that the propagator isn’t invariant under all symplectomorphisms, as is assumed.
Remark 4.1.3.
The map is always asymptotically constant as ([28]), and integrates a perturbed connection — the curvature of this perturbed connection is the Bergman Kähler form, which is an -perturbation of the original symplectic form, [26], [33]. Therefore, it’s still true that
which is what is needed for a quantization. See remark 2.0.2. When is constant we have an exact equality.242424The property of the Bergman kernel being constant was used in the context of Kähler quantization in [4]. For it to be constant, it is enough that the manifold and line bundle are homogeneous.
The function is a well-studied map, and it being constant has other interesting consequences, eg. it implies stability in GIT (see page 2 of [27]). Metrics such that this map is constant are called balanced, [8]. For a generalization of the Bergman kernel to symplectic manifolds, see section 4 of [29] and [15].
4.2 Surfaces of Constant Curvature
Here we will recall some explicit examples of the coherent state path integral eq. 0.0.1 computed by Daubechies, Klauder, on Kähler manifolds in [7], [16]. We will show that they satisfy definition 2.0.1. To arrive at these expressions, the aforementioned authors (essentially) introduce Brownian motion into eq. 0.0.1 via a massive term and send 252525For more explicit examples of Bergman kernels, see [15]. In all of the following, the kernels are constant along the diagonal. In this paper, we have argued that this is true in general: whenever the Bergman kernel is constant along the diagonal, it computes the coherent state path integral.
Example 4.2.1.
The following locally describe some 2-dimensional Kähler manifolds and their propagators eq. 0.0.1 in Darboux coordinates, for which the connection on the prequantum line bundle is and is the Riemannian metric (which has constant curvature).
1.
Flat space:
(4.2.1) (4.2.2) -
2.
Sphere:
(4.2.3) (4.2.4) -
3.
Hyperbolic plane: 262626The examples computed by Klauder depend on parameters which we’ve set equal to
(4.2.5) (4.2.6)
A simple computation shows that, indeed,
(where the exterior derivative is taken in Therefore, integrates in the sense of definition 2.0.1. This illustrates the non-uniqueness of the value of eq. 0.0.1, since the complement of an arc connecting the north and south pole of the sphere (which has measure zero) is symplectomorphic to with its standard symplectic form.
Let us emphasize the following: there is a unique solution to eq. 0.0.5, 0.0.6, while the corresponding solution to eq. 1.1.2-1.1.5 (or eq. 2.0.4-2.0.6) is not unique, and this is for the same reason that the Riemann integral doesn’t have a unique value under path integrals, as discussed in section 3.2 — in order to produce a unique value, one needs to specify a second order term, as in eq. 3.2.5. This leads us into the following:
4.2.1 The Second Order Taylor Expansion of the Propagator (Bergman Kernel)
Since the first order Taylor expansion doesn’t uniquely determine the propagator, more interesting then is the second order Taylor expansion in the variables of at This does distinguish between the different propagators. To make sense of this, we use the corresponding Riemannian metric to split the short exact sequence associated to the jet bundles
We do this because, in order to define higher order Taylor expansions on manifolds, we need to split this sequence. Given a Riemannian metric, the result is the usual Taylor expansion formula, with derivatives replaced by symmetrized covariant derivatives, and terms like replaced by
For the hyperbolic plane, the non-zero Christoffel symbols are given by
Proposition 4.2.2.
In all three cases, the second order Taylor expansions are given by
Geometrically, this is equal to
where we identify with the corresponding vertical vector induced by the -action.
Classically (ie. over smooth paths ), the higher order term doesn’t contribute to the integral (see section 3), ie.
Therefore, it is more precise to say that these propagators are equal to
with the understanding that eq. 4.2.12 doesn’t hold on non-smooth paths. It would be interesting to know if the second order Taylor expansions (or 2-jets) uniquely determine the propagator — these 2-jets are given by a higher order van Est map ([22]).
4.2.2 The Lattice Approach to the Path Integral
Similarly, it would be interesting to know which path integral approximations given by the lattice approach of [22] converge to the propagator. This involves choosing a section of over a neighborhood of the diagonal in such that: restricted to the diagonal is equal to and its derivative in the second component, evaluated on the diagonal, agrees with the connection — this means that the connection determined by agrees with 272727We can always use the the Bergman kernel itself for , in which case the following lattice construction clearly works.
For example, we could cover with small enough open sets such that there is a unique geodesic connecting any two points in any of these open sets — this determines a neighborhood of the diagonal in with if for some We get a section of over with the correct derivative, given by
where is the distance between the corresponding points and denotes parallel transport over the geodesic connecting to For a map and for a triangulation of we get an approximation to the integrand of eq. 4.2.13 given by
We can then compute the path integral eq. 4.2.13 as in Feynman’s approach, where the measure is approximated using and the normalization is determined by the fact that the restriction of the path integral to the diagonal is equal to the That is, we can try to compute
where and is determined by the fact that for each the approximation is equal to if
Note that indeed, the derivative (or pushforward) of in the second component, evaluated at the diagonal, agrees with Classicaly, this is enough information to get a well-defined integral (see section 3) — the point is that, this is all of the information that is determined by the formal path integral. We can use the metric to split the short exact sequence associated to the jet bundle, eq. 4.2.8, and using this splitting also has the correct second derivative per eq. 4.2.13, ie.
5 Relation to Other Quantization Schemes
5.1 Relation to Kontsevich’s Work, the Poisson Sigma Model and Weinstein’s Program
In the context of symplectic manifolds, the propagator is a section of the prequantum line bundle over the symplectic groupoid, and it is an idempotent with respect to the twisted convolution algebra ([13]). Therefore, definition 2.0.1 naturally generalizes to Poisson manifolds. Equation 2.0.5 implies that
where is the Poisson bivector and
is a function defined on composable arrows. To understand this function, consider three marked points on Equation 5.1.2 should be interpreted as
where the integral is over morphisms whose restrictions to
are algebroid-homotopic to respectively, and where is any morphism whose restriction to is 303030The symplectic groupoid being prequantizable implies that for morphisms hence this is well-defined. Therefore, on a Poisson manifold this construction gives a -algebra and quantization map, as in Weinstein’s program [36], and can be used to compute
Here, the integral is over all contractible morphisms and is any morphism whose restriction to is This is formally equivalent to the Poisson sigma model approach to Kontsevich’s star product [19], see [2]. In the symplectic case, it is the same as eq. 0.0.2 (after gauge reduction). See section 1.4.2.
5.2 Relation to Kostant-Souriau’s Geometric Quantization
In section 1.3 we discussed the relation of these propagators to geometric quantization, using Kähler polarizations (also see section 5.1). However, one can ask about real polarizations.313131Of course, there are examples where the propagator can be indirectly obtained using real polarizations, as discussed in section 1.3. In this section we discuss viewing real-polarized sections as the physical states associated to a propagator. In [24] (also see [23]) we discussed a function
satisfying all of the conditions to be a propagator integrating except that rather than satisfying it satisfies the relaxed condition
Therefore, it’s not a projection, but it still leads to good quantizations. In particular, the corresponding eigenspace consists of all (complex)-linearly polarized sections (ie. for all linear polarizations), and on this subspace does give the correct inner product — the fact that all (complex) linearly-polarized sections are in this eigenspace leads to natural unitary equivalences between all such quantizations. Furthermore, it leads to a non-perturbative deformation quantization, which simultaneously acts on all linearly-polarized sections.
It seems that we can view the failure of as a failure of Fubini’s theorem, since
is not in for fixed This can still be considered to be a propagator, just not in the precise sense defined in this paper — though it seems this generalization is useful. It appears that any quantization scheme is (implicitly)
attempting to assign a value to eq. 0.0.1, though traditional geometric quantization forgets a lot of information about the quantization map.
Question: When does there exist an integrating such that
for some
.3 The van Est map and Riemann Sums on Manifolds
The natural setting for the construction of the (generalized) Riemann sums on manifolds we give is the category of Lie groupoids, but it only involves the simplest Lie groupoid that exists, the pair groupoid. The pair groupoid of a manifold is the groupoid with a unique arrow between any two points in This is discussed more in [21], [22].
Definition .3.1.
323232This formulation is presented in [22], section 3. Also, [21]. It is equivalent to, but a bit different from, the original formulation of Weinstein-Xu in [37]. It was generalized (to eg. -valued functions) in [20], which is relevant to path integrals.Let be a manifold. Let be a smooth function which vanishes on the diagonal and which is invariant under even permutations. The van Est map (in degree 0) applied to gives an -form on Letting be vectors at a point it is defined by
where differentiates in the component (where occupies the component).
Example .3.2.
Let and let be a function on which satisfies the conditions of the previous definition. Then
States differently, is a normalized -cochain on which is invariant under even permutations.
Now, let be an oriented, -dimensional manifold. Given a triangulation of up to even permutation there is a canonical ordering of the vertices of each –dimensional simplex. Therefore, we can evaluate on any -dimensional face by choosing such an ordering of the vertices and plugging them into This is well-defined because we are assuming that is invariant under even permutations. We denote its value by
Definition .3.3.
Theorem .3.4.
([21], [22]) Suppose that satisfies the conditions of definition .3.1 and that Then
where the limit is taken over barycentric subdivisions of any triangulation
The next proposition shows that this notion of Riemann sum is well–behaved with respect to pullbacks, which is important because pullbacks appear in path integrals:
Proposition .3.5.
Let be a triangulation of an -dimensional manifold let be a smooth function and let be an -form on with Then
is a (generalized) Riemann sum of where is the pullback via the map induced by